Question 33 5 pts If the required-reserve ratio is 10 percent and the Fed buys $25 million worth of government bonds, the maximum potential change in the money supply will be a(n): increase of $250 million. increase of $25 million. decrease of $25 million. decrease of $250 million.
Answer #1

Money multiplier = 1/RR = 1/0.10 = 10

The Fed buys $25 million worth of government bonds , the maximum potential change in the money supply will be an increase of (10)($25 million) = $250 million. Hence,option(A) is correct.